The probability of a customer arrival at a grocery service counter in any 1 second is equal to 0,1. Assume that customers arrive in a random stream and hence that the arrival any 1 second is independent of any other. Calculate the probability that the first arrival will occur during the third   1-second interval.


1
Expert's answer
2021-06-15T13:51:18-0400

"p=0.1 \\

q=1-q=0.9"

Geometric probability:

"p(y) = q^{y-1}p \\

p(3) = 0.9^{3-1} \times 0.1 = 0.081 = 8.1 \;\%"